# A non-singular universe out of Hayward black hole

### v2

@@ Abstract
We construct a (quantum mechanically) modified model for the Oppenheimer-Snyder collapse scenario where the exterior of the collapsing dust ball is a Hayward black hole spacetime and the interior is a dust Friedmann-Robertson-Walker cosmology. This interior cosmology is entirely determined by the junction conditions with the exterior black hole. It turns out to be non-singular, displaying a power-law contraction which precedes a de Sitter phase or, reversely, a power-law expansion followed by a de Sitter era. We also analyse the global causal structure and the viability of the model.
@@ I Introduction
The question of whether the spacetime singularities inherent to General Relativity (GR) represent a real physical phenomenon remains open more than a century after the advent of this gravitational theory. In cosmological scenarios, the debate affects fundamental questions about the origin of the Universe. Was there something before the Big Bang? Did the Universe have a beginning? In black hole physics, the problem is related to other important physical issues. What is the fate of free-falling observers? What happens beyond the event horizon? The classical Oppenheimer-Snyder (OS) collapse scenario [1] is a especially suitable arena to discuss the connection between the aforementioned questions. The fact that we can model a collapsing dust ball using a Friedmann-Robertson-Walker (FRW) metric in its interior and a Schwarzschild solution in the exterior suggests a link between black hole spacetimes and cosmology. Nevertheless, the resulting geometry still possesses a singularity, posing serious difficulties at the classical level and, more importantly, sustaining the black hole information paradox in a semiclassical picture of the black hole evaporation [2].
@@
The singular nature of the Schwarzschild spacetime has motivated considerable effort devoted to exploring alternatives [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. A prominent example is the Hayward black hole [5], on which we will focus our attention in this work. Instead of the singularity at the origin of the radial coordinate, this spacetime displays a regular ‘quantum gravity core’. The Hayward geometry has a really special conformal diagram, representing an evaporating non-singular black hole (in which back-reaction is effectively introduced) [5]111 Although the associated static metric was originally derived by Poisson & Israel from the model inspired by one-loop corrections to Einstein field equations [3], its relevance for the information puzzle was established in Ref. [5].. All matter entering the black hole ultimately reaches future null infinity. Hence, in a semiclassical approximation, most probably there is enough Cauchy data at this future infinity to reconstruct the past, resolving the information paradox.
@@
Although the Hayward metric seems to solve some fundamental problems present in the Schwarzchild metric, there has been no consensus about which physical theory the metric could be derived from. However, recently the metric has been derived as a vacuum solution within two independent effective descriptions incorporating quantum gravity corrections [39, 29]. The first one, called Quasi-Topological Theory [40, 39], can be regarded as Einstein-Hilbert action supplemented with an infinite tower of higher-curvature corrections which are contractions of the Riemann tensor. For instance, when the tower is truncated at the second order, one gets Gauss-Bonnet gravity (see Refs. [41, 42]). In general, such an infinite tower of corrections, weighted by powers of e.g. the Planck length, is what one could expect from String Theory (ST) as an ultraviolet completion of GR [42]. Specifically, when the number of spacetime dimensions $D$ is equal or greater than five and spherical symmetry is imposed, the equations of motions (being of second differential order) determine the Hayward metric as a unique vacuum solution, provided that the simplest sequence of gravitational couplings is chosen [39]. The other aforementioned description [29], built in four spacetime dimensions, leads to the Hayward metric as the vacuum solution coming from a deformed (or “polymerized”) gravitational Hamiltonian in Lemaître-Tolman-Bondi (LTB) spacetimes. Motivated by dust collapse models in Loop Quantum Gravity (LQG), where the effective dynamics can be regarded as an LTB model with a polymerized Hamiltonian [27, 28, 29], the authors of Ref. [29] construct a specific “polymerization function” for the Hamiltonian that reproduces the Hayward metric in vacuo. Contrary to what happens in standard effective LQG models, the polymerization function is now unbounded. Moreover, this description can be translated into a Lagrangian mimetic gravity theory with a concrete mimetic potential.
@@
In the present work, we will assume that the Hayward metric describes the vacuum exterior of a modified OS collapse scenario. We will then impose conditions for the interior compatible with the above descriptions and derive the corresponding cosmological behavior. Indeed, we will explore the resulting interior dust cosmology when we require: a) both the exterior and interior metrics join smoothly at the boundary, located at the surface of the collapsing dust ball, b) a conservation law for dust is satisfied. This investigation should open the door to exploring how well the interior FRW background describes our universe.
@@
The interior geometry will be completely determined just by imposing Israel-Darmois junction conditions at the boundary. The mathematical implementation of this condition and the formalism on which it is based were already studied by Pawłowski and the author in Refs. [23, 24], in the context of some effective models of OS collapse in LQG. In those previous works, it was the exterior metric that was determined in terms of the bouncing cosmology. More recently, similar techniques were also applied by several other authors [25, 43, 28, 27, 29, 36, 44]. The dust cosmology that we will obtain has the following remarkable features. It has no curvature singularities, it presents a de Sitter era preceding (following) a power-law expansion (contraction), and the transition between these two regimes is smooth, displaying a graceful exit. Moreover, we will see that our modified OS model shares the behavior found in Ref. [44] for a model inspired by Asymptotically Safe Gravity (ASG). Finally, we will also construct and discuss the conformal diagram representing the modified OS collapse (without back-reaction due to Hawking radiation).
@@ II Cosmological dynamics from junction conditions
where $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\varphi^{2}$ is the metric on the two-sphere, and the function $F(X)$ is chosen to correspond to a Hayward black hole [5], namely
@@
Here, $M$ is a ‘mass’ constant. For large values of the radial coordinate $X$ the metric behaves like the Schwarzchild one. The other constant in Eq. (2) is usually taken to be the Planck length (and hence $l=1$ in the Planck units that we are adopting), although we will maintain it explicitly in our formulae. For small $X$, we have $F(X)\sim 1-X^{2}/l^{2}$. Thus, we can expect that the cosmological interior that matches smoothly at the boundary displays an epoch of de Sitter behavior. Throughout the rest of this article, we restrict the analysis to the non-extreme case when $M>(3\sqrt{3}/4)l$ in Eq. (2), so that the mass ranges over values of astrophysical interest. In this case, $F(X)$ has two positive roots, say $0<X^{-}<X^{+}$.
@@
The form of the scale factor $a(T)$ is determined by our junction conditions with the exterior. In full detail, we demand that the Israel-Darmois junction conditions (see e.g. Chapters 3.7-3.8 of Ref. [45]) be satisfied at the surface of the collapsing dust ball. We call $\Sigma$ this boundary. Thus, the induced metric on $\Sigma$ and its extrinsic curvature (second fundamental form) must coincide when computed from the exterior and from the interior, namely, $h_{ab}^{\mathrm{ext}}=h_{ab}^{\mathrm{int}}$ and $K_{ab}^{\mathrm{ext}}=K_{ab}^{\mathrm{int}}$, respectively. A detailed analysis of these conditions on the line elements (1) and (3) was recently carried out in Refs. [24, 25, 46] [^2] See Ref. [46] for the general analysis of the case where the exterior line element (1) is not necessarily stationary.. Here, we will summarize the most important aspects. Let $T$ be a parameter at $\Sigma$. The trajectory of the free-falling observer at $\Sigma$ can be described either as $x^{\alpha}_{\mathrm{ext}}=(t=\tau(T),X=R(T)$) from the exterior or as $x^{\alpha}_{\mathrm{int}}=(T,r=r_{b}$) from the interior, where $r_{b}$ is a constant. The junction conditions form a soluble system of algebraic equations. They imply the following relations [24, 25, 46], that are crucial for this work:
@@
It is challenging to invert this equation to get $a(T)$. However, we can analyse it to deduce important properties of the resulting scale factor. One has $\lim_{a\rightarrow\infty}T(a)=-\infty$ and $\lim_{a\rightarrow 0}T(a)=\infty$. These limits, together with the fact that relation (7) is monotonic in $a\in(0,\infty)$, indicate that the inverse function $a(T)$ describes a timelike geodesically complete dust universe, in which matter is contracting for all values of the affine parameter $T\in(-\infty,\infty)$. Thanks to homogeneity, the scale factor obtained by solving the junction conditions is valid not only at $\Sigma$, but in the whole interior. Note that the $\log$ term dominates in Eq. (7) for small $a$, implying that the contraction of the Universe is exponential in this regime. On the other hand, for large $a$, the $\log$ contribution becomes negligible and the scale factor approximately behaves as for dust collapse in GR, namely, $a(T)\sim T^{2/3}$. Since the dependence of $T(a)$ given in Eq. (7) is smooth and strictly monotonic for $a\in(0,\infty)$, the deduced interior geometry admits a graceful exit from the de Sitter epoch. It is worth remarking that the Krestchmann scalar $K_{\mathrm{int}}$ is bounded everywhere in the interior (see Appendix A). In particular, in the limit $a\rightarrow 0$, when the metric (3) becomes degenerate, we have $K_{\mathrm{int}}\rightarrow 24/l^{4}$.
@@
where $H=\frac{\dot{a}}{a}$ and the prime stands for the derivative with respect to $a$. Condition (8) is the consequence of the conservation of the energy-momentum tensor for dust, $\nabla_{\mu}\mathcal{T}^{\mu\nu}=0$, where $\mathcal{T}^{\mu\nu}=\rho k^{\mu}k^{\nu}$ and $k^{\mu}\partial_{\mu}=\partial_{T}$. It is compatible with the two mentioned theories that identify the Hayward metric as a vacuum solution [39, 41, 29] [^3] The conservation law indeed holds for the four-dimensional variant of the theory with an infinite tower of higher-order corrections when the FRW ansatz is adopted (see the discussion around Eq. (7) in Ref. [41]). . The corrected Friedmann equations given above reproduce the dynamical equations of a dust FRW cosmology in GR both in the low energy limit (at leading order when $\rho$ is small) and in the limit $l\rightarrow 0$. Note also that, in the limit $a\rightarrow 0$, the energy density is clearly infinite. This issue will be discussed in Sec. IV.

### v3

# Cosmic inflation prevents singularity formation in collapse into a Hayward black hole

@@ Abstract
We construct a (quantum mechanically) modified model for the Oppenheimer-Snyder collapse scenario where the exterior of the collapsing dust ball is a Hayward black hole spacetime and the interior is a dust Friedmann-Robertson-Walker cosmology. This interior cosmology is entirely determined by the junction conditions with the exterior black hole. It turns out to be non-singular, displaying a power-law contraction which precedes a de Sitter phase or, reversely, a power-law expansion followed by a de Sitter era. We demonstrate that cosmic inflation in the collapse setting is a mechanism that decelerates collapsing matter, thereby preventing singularity formation. We also analyse the global causal structure and the viability of the model.
@@ I Introduction
The question of whether the spacetime singularities inherent to General Relativity (GR) represent a real physical phenomenon remains open more than a century after the advent of this gravitational theory. In cosmological scenarios, the debate affects fundamental questions about the origin of the Universe. Was there something before the Big Bang? Did the Universe have a beginning? In black hole physics, the problem is related to other important physical issues. What is the fate of free-falling observers? What happens beyond the event horizon? The classical Oppenheimer-Snyder (OS) collapse scenario Oppenheimer and Snyder (1939) is a especially suitable arena to discuss the connection between the aforementioned questions. The fact that we can model a collapsing dust ball using a Friedmann-Robertson-Walker (FRW) metric in its interior and a Schwarzschild solution in the exterior suggests a link between black hole spacetimes and cosmology. Nevertheless, the resulting geometry still possesses a singularity, posing serious difficulties at the classical level and, more importantly, sustaining the black hole information paradox in a semiclassical picture of the black hole evaporation Ashtekar and Bojowald (2005).
@@
The singular nature of the Schwarzschild spacetime has motivated considerable effort devoted to exploring alternatives Poisson and Israel (1988); Dymnikova (1992); Hayward (2006); Maier and Soares (2009); Barragan et al. (2009); Bambi et al. (2013); Saini and Stojkovic (2014); De Lorenzo et al. (2015); Frolov (2015); Frolov et al. (2015); Frolov (2016); Carballo-Rubio et al. (2018); Kiefer and Schmitz (2019); Mosani et al. (2020); Schmitz (2021); Góźdź et al. (2023); Bonanno et al. (2023); Shafiee and Bahrampour (2023); Kiefer and Mohaddes (2023); Shojai et al. (2022); Malafarina and Toshmatov (2022); Alonso-Bardaji and Brizuela (2024); Duque (2024); Barca and Montani (2024); Bobula ; Bobula and Pawłowski (2023); Lewandowski et al. (2023); Giesel et al. (2023, 2024a, 2024b, 2024c); Kelly et al. (2021, 2020); Husain et al. (2022a, b); Fazzini et al. (2024, 2023); Cafaro and Lewandowski (2024); Cipriani et al. (2024); Han et al. (2024). A prominent example is the Hayward black hole Hayward (2006), on which we will focus our attention in this work. Instead of the singularity at the origin of the radial coordinate, this spacetime displays a regular ‘quantum gravity core’. The Hayward geometry has a really special conformal diagram, representing an evaporating non-singular black hole (in which back-reaction is effectively introduced) Hayward (2006)111 Although the associated static metric was originally derived by Poisson & Israel from the model inspired by one-loop corrections to Einstein field equations Poisson and Israel (1988), its relevance for the information puzzle was established in Ref. Hayward (2006).. All matter entering the black hole ultimately reaches future null infinity. Hence, in a semiclassical approximation, most probably there is enough Cauchy data at this future infinity to reconstruct the past, resolving the information paradox.
@@
Although the Hayward metric seems to solve some fundamental problems present in the Schwarzchild metric, there has been no consensus about which physical theory the metric could be derived from. However, recently the metric has been derived as a vacuum solution within two independent effective descriptions incorporating quantum gravity corrections Bueno et al. (2025); Giesel et al. (2024c). The first one, called Quasi-Topological Theory Aguilar-Gutierrez et al. (2023); Bueno et al. (2025), can be regarded as Einstein-Hilbert action supplemented with an infinite tower of higher-curvature corrections which are contractions of the Riemann tensor. For instance, when the tower is truncated at the second order, one gets Gauss-Bonnet gravity (see Refs. Arciniega et al. (2020); Bueno et al. (2024)). In general, such an infinite tower of corrections, weighted by powers of e.g. the Planck length, is what one could expect from String Theory (ST) as an ultraviolet completion of GR Bueno et al. (2024). Specifically, when the number of spacetime dimensions $D$ is equal or greater than five and spherical symmetry is imposed, the equations of motions (being of second differential order) determine the Hayward metric as a unique vacuum solution, provided that the simplest sequence of gravitational couplings is chosen Bueno et al. (2025). The other aforementioned description Giesel et al. (2024c), built in four spacetime dimensions, leads to the Hayward metric as the vacuum solution coming from a deformed (or “polymerized”) gravitational Hamiltonian in Lemaître-Tolman-Bondi (LTB) spacetimes. Motivated by dust collapse models in Loop Quantum Gravity (LQG), where the effective dynamics can be regarded as an LTB model with a polymerized Hamiltonian Giesel et al. (2024a, b, c), the authors of Ref. Giesel et al. (2024c) construct a specific “polymerization function” for the Hamiltonian that reproduces the Hayward metric in vacuo. Contrary to what happens in standard effective LQG models, the polymerization function is now unbounded. Moreover, this description can be translated into a Lagrangian mimetic gravity theory with a concrete mimetic potential.
@@
In the present work, we will assume that the Hayward metric describes the vacuum exterior of a modified OS collapse scenario. We will then impose conditions for the interior that we expect are compatible with the above descriptions and derive the corresponding cosmological behavior. Indeed, we will explore the resulting interior dust cosmology when we require: a) both the exterior and interior metrics join smoothly at the boundary, located at the surface of the collapsing dust ball, b) a conservation law for dust is satisfied. This investigation should open the door to exploring how well the interior FRW background describes our universe.
The interior geometry will be completely determined by imposing the Israel-Darmois junction conditions at the boundary. Within GR, Israel-Darmois junction conditions require induced metric and extrinsic curvature to be continuous at the junction surface provided no energy layer or thin shell being present there, which is the case for OS collapse scenario. We point out that these junction conditions were originally derived for General Relativity, thus, in general, when working with dynamical equations different from the Einstein Field Equations (EFEs), other conditions might be required Chu and Tan (2022). However, based purely on geometric considerations — without reference to any dynamical equations — we observe that the Gauss-Codazzi equations (see Eqs. 3.39–3.40 in Ref. Poisson (2009)) relate the extrinsic curvature to the full Riemann tensor evaluated on a hypersurface (the junction surface). Consequently, if we desire the full Riemann tensor to be continuous at the junction surface, the Gauss-Codazzi equations suggest that the extrinsic curvature must also be continuous on that surface (and vice versa). Therefore, we expect (but we do not prove) this geometric requirement — namely, the continuity of the extrinsic curvature — to hold in frameworks that identify the Hayward black hole as the unique vacuum solution in the context of Oppenheimer-Snyder collapse. In these descriptions, the connection is Levi-Civita, and the aforementioned Gauss-Codazzi equations remain valid. Indeed, since these frameworks do not introduce additional matter fields or degrees of freedom beyond those in GR, we expect that the requirement of metric and extrinsic curvature continuity at the surface of the dust ball (junction surface) is compatible with them. As we will show, merely requiring continuities of metric and extrinsic curvatures (Israel-Darmois junction conditions) fully determines the interior FRW geometry. However, we do not rigorously derive junction conditions viable for Quasi-Topological gravity and effective LQG-inspired frameworks. As shown in Ref. Chu and Tan (2022) for certain theories beyond GR, imposing Israel-Darmois junction conditions is inconsistent with a vanishing surface energy-momentum tensor or additional, nontrivial relations between the extrinsic and intrinsic curvatures have to be satisfied. For these reasons, we emphasize that the validity of the Israel-Darmois junction conditions remain an essential assumption of our work. To further support the application of the Israel-Darmois conditions we observe that i) in Gauss-Bonnet gravity (which can be regarded as Quasi-Topological gravity truncated at second order), the requirement of continuity of extrinsic curvature is viable (see, e.g., the discussion around Eqs. 3.67–3.68 in Ref. Chu and Tan (2022)) and ii) in the framework of LQG-inspired LTB models (however, with the standard polymerization function), it was demonstrated in Ref. [40] that applying the Israel-Darmois junction conditions and solving the dynamical equations yield the same results in the modified Oppenheimer-Snyder collapse scenario. As an aside, note that the formalism of Israel-Darmois junction conditions in the context of Oppenheimer-Snyder collapse has been successfully applied to models beyond the Einstein Field Equations by several authors in Refs. Maier and Soares (2009); Schmitz (2021); Malafarina and Toshmatov (2022); Bobula ; Lewandowski et al. (2023); Bobula and Pawłowski (2023); Han et al. (2023); Giesel et al. (2024b, a, c); Cafaro and Lewandowski (2024); Bonanno et al. (2024).
@@
The dust cosmology that we will obtain has the following remarkable features. It has no curvature singularities, it presents a de Sitter era preceding (following) a power-law expansion (contraction), and the transition between these two regimes is smooth, displaying a graceful exit. Moreover, we will see that our modified OS model shares the behavior found in Ref. Bonanno et al. (2024) for a model inspired by Asymptotically Safe Gravity (ASG). Finally, we will also construct and discuss the conformal diagram representing the modified OS collapse (without back-reaction due to Hawking radiation).
@@ II Cosmological dynamics from junction conditions
where $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\varphi^{2}$ is the metric on the two-sphere, and the function $F(X)$ is chosen to correspond to a Hayward black hole Hayward (2006), namely
@@
Here, $M$ is a ‘mass’ constant. For large values of the radial coordinate $X$ the metric behaves like the Schwarzchild one. The other constant in Eq. (2) is usually taken to be the Planck length (and hence $l=1$ in the Planck units that we are adopting), although we will maintain it explicitly in our formulae. For small $X$, we have $F(X)\sim 1-X^{2}/l^{2}$. Thus, we can expect that the cosmological interior that matches smoothly at the boundary displays an epoch of de Sitter behavior. Throughout the rest of this article, we restrict the analysis to the non-extreme case when $M>(3\sqrt{3}/4)l$ in Eq. (2), so that the mass ranges over values of astrophysical interest. In this case, $F(X)$ has two positive roots, say $0<X^{-}<X^{+}$. For the remaining cases $M\leq(3\sqrt{3}/4)l$ the causal structures will be different, however, the corresponding analysis is beyond the scope of this work.
@@
The form of the scale factor $a(T)$ is determined by our junction conditions with the exterior. In full detail, we demand that the Israel-Darmois junction conditions (see e.g. Chapters 3.7-3.8 of Ref. Poisson (2009)) be satisfied at the surface of the collapsing dust ball. We call $\Sigma$ this boundary. Thus, the induced metric on $\Sigma$ and its extrinsic curvature (second fundamental form) must coincide when computed from the exterior and from the interior, namely, $h_{ab}^{\mathrm{ext}}=h_{ab}^{\mathrm{int}}$ and $K_{ab}^{\mathrm{ext}}=K_{ab}^{\mathrm{int}}$, respectively. A detailed analysis of these conditions on the line elements (1) and (3) was recently carried out in Refs. Bobula and Pawłowski (2023); Lewandowski et al. (2023); Bobula and Pawłowski (2024) [^2] See Ref. Bobula and Pawłowski (2024) for the general analysis of the case where the exterior line element (1) is not necessarily stationary.. Here, we will summarize the most important aspects. Let $T$ be a parameter at $\Sigma$. The trajectory of the free-falling observer at $\Sigma$ can be described either as $x^{\alpha}_{\mathrm{ext}}=(t=\tau(T),X=R(T)$) from the exterior or as $x^{\alpha}_{\mathrm{int}}=(T,r=r_{b}$) from the interior, where $r_{b}$ is a constant. The junction conditions form a soluble system of algebraic equations. They imply the following relations Bobula and Pawłowski (2023); Lewandowski et al. (2023); Bobula and Pawłowski (2024), that are crucial for this work:
@@
It is challenging to invert this equation to get $a(T)$. However, we can analyse it to deduce important properties of the resulting scale factor. One has $\lim_{a\rightarrow\infty}T(a)=-\infty$ and $\lim_{a\rightarrow 0}T(a)=\infty$. These limits, together with the fact that relation (7) is monotonic in $a\in(0,\infty)$, indicate that the inverse function $a(T)$ describes a timelike geodesically complete dust universe, in which matter is contracting for all values of the affine parameter $T\in(-\infty,\infty)$. Thanks to homogeneity, the scale factor obtained by solving the junction conditions is valid not only at $\Sigma$, but in the whole interior. Note that the $\log$ term dominates in Eq. (7) for small $a$, implying that the contraction of the Universe is exponential in this regime. On the other hand, for large $a$, the $\log$ contribution becomes negligible and the scale factor approximately behaves as for dust collapse in GR, namely, $a(T)\sim T^{2/3}$. Since the dependence of $T(a)$ given in Eq. (7) is smooth and strictly monotonic for $a\in(0,\infty)$, the deduced interior geometry admits a graceful entrance to the de Sitter epoch. It is worth remarking that the Kretschmann scalar $\mathcal{K}_{\mathrm{int}}$ is bounded everywhere in the interior. Indeed, this curvature invariant for the line element (3), with the scale factor $a$ determined by the inverse of $T(a)$ given in Eq. (7), can be calculated to be
$\begin{split}\mathcal{K}_{\mathrm{int}}&=\frac{12\left(a^{2}T^{\prime\prime}(a)^{2}+T^{\prime}(a)^{2}\right)}{a^{4}T^{\prime}(a)^{6}}=\\ &=\frac{12M^{2}\left(5a^{6}r_{b}^{6}+8a^{3}l^{2}Mr_{b}^{3}+32l^{4}M^{2}\right)}{\left(a^{3}r_{b}^{3}+2l^{2}M\right)^{4}}\,.\end{split}$ (8)
In particular, in the limit $a\rightarrow 0$, when the metric (3) becomes degenerate, we have $\mathcal{K}_{\mathrm{int}}\rightarrow 24/l^{4}$.
@@
where $H=\frac{\dot{a}}{a}$ and the prime stands for the derivative with respect to $a$. Condition (9) is the consequence of the conservation of the energy-momentum tensor for dust, $\nabla_{\mu}\mathcal{T}^{\mu\nu}=0$, where $\mathcal{T}^{\mu\nu}=\rho k^{\mu}k^{\nu}$ and $k^{\mu}\partial_{\mu}=\partial_{T}$. It is compatible with the two mentioned theories that identify the Hayward metric as a vacuum solution Bueno et al. (2025); Arciniega et al. (2020); Giesel et al. (2024c) [^3] The conservation law indeed holds for the four-dimensional variant of the theory with an infinite tower of higher-order corrections when the FRW ansatz is adopted (see the discussion around Eq. (7) in Ref. Arciniega et al. (2020)). . The corrected Friedmann equations given above reproduce the dynamical equations of a dust FRW cosmology in GR both in the low energy limit (at leading order when $\rho$ is small) and in the limit $l\rightarrow 0$. Note also that, in the limit $a\rightarrow 0$, the energy density is clearly infinite. This issue will be discussed in Sec. IV.
Figure 2: Left: the conformal diagram representing our (quantum mechanically) modified OS collapse scenario. The FRW geometry, enclosed by curves marked by $r=0$ and $r=r_{b}$, is surrounded by the static Hayward black hole geometry, with horizons $X=X_{-}$ and $X=X_{+}$. Right: the conformal diagram for the classical OS collapse scenarioOppenheimer and Snyder (1939). Singular FRW geometry, whose dynamics is governed by the classical dust Friedmann equation $\left(\frac{\dot{a}}{a}\right)^{2}=\frac{8\pi\rho}{3}$ is surrounded by Schwarzschild black hole geometry — the metric function for the exterior is $F\rightarrow F_{\mathrm{Sch}}(X)=1-\frac{2M}{X}$ with Schwarzchild radius located at $X=X_{\mathrm{Schw}}=2M$ and the singularity $X=0$. To generate both of the diagrams we have taken the same constants’ values $M=15$, $a_{0}=5$, $r_{b}=(4\pi/3)^{-1/3}$ and $\kappa=0.1$ (in Planck units).